OpenStudy (anonymous):

What is the value of the fourth term in a geometric sequence for which a1=10 and r=0.5

OpenStudy (anonymous):

a2 = r a1 a3 = r a2 = r^2 a1 a4 = r a3 = r^3 a1 plug in a1 and r for answer

OpenStudy (anonymous):

Use the general term of G.P. an=ar^n-1

OpenStudy (anonymous):

APEX 1.25